### v1

@@ Abstract
We present an interactive visualization system for exploring named entities and their relationships across document collections. The system is designed around a graph-based representation that integrates three types of nodes: documents, entity mentions, and entities. Connections capture two key relationship types: (i) identical entities across contexts, and (ii) co-locations of mentions within documents. Multiple coordinated views enable users to examine entity occurrences, discover clusters of related mentions, and explore higher-level entity group relationships. To support flexible and iterative exploration, the interface offers fuzzy views with approximate connections, as well as tools for interactively editing the graph by adding or removing links, entities, and mentions, as well as editing entity terms. Additional interaction features include filtering, mini-map navigation, and export options to JSON or image formats for downstream analysis and reporting. This approach contributes to human-centered exploration of entity-rich text data by combining graph visualization, interactive refinement, and adaptable perspectives on relationships.
@@ 4 System Design
NERVIS is an interactive web-based system for graph visualization and editing, implemented using HTML, CSS, TypeScript, and React framework [^1] https://react.dev. The system is designed around a modular architecture that separates data management, visualization, and user interaction. At its core, NERVIS maintains structured data models that represent nodes, edges, and associated metadata, enabling efficient storage, retrieval, and manipulation of graph information. These models feed into a graph visualization pipeline, which computes layouts, applies styling and interaction rules, and renders the graph for exploration and analysis.
@@ 4.2 Visualization Pipeline
Figure 2: NERVIS visualization pipeline. Orange and gey nodes indicate intractable and uninterpretable stages respectively. Green nodes indicate intermediary data representations.
@@ 5 Expert Evaluation
After initial consultation set the stage and expectations, we performed two rounds of expert system evaluation with synthetic data. After preliminary evaluation, the feature requests were prioritized and the most important ones were implemented. After the final round, the expert judged the system to be mature enough for use in real-life setting. They also composed the list of possible future improvements, additions and extensions.
In the initial consultation we discussed software goals and previous experience when handling named-entity data with existing software. One of the key things pointed out during this consultation was the burden of human labor when dealing with computer-generated labels. For historians, a recall-focused approach is preferred to an accuracy-focused one. They prefer broad sets of result candidates, which they can browse through and edit manually as opposed to only seeing a single most likely result. This encouraged the design choice to develop a tool where the deletion of nodes and edges is the most common user operation, as opposed to creating new ones. A high recall approach inevitably generates relatively high amounts of noise, which we planned to tackle with a variety of content filters that allow researchers to focus on specific documents, entities, or types when cleaning the data. Since the users of Digital Humanities tools often do not have an engineering background and often struggle with the installation process, we decided on the web platform, which also makes the system broadly available.
After the prototype was tested, we followed up with our second consultation. At this stage, the tool received generally positive feedback, with feature requests focusing on the most common user actions. As already mentioned, with high recall results, most user actions involve deleting nodes and edges, which was made additionally accessible with the implementation of hotkeys for quick node/edge deletion. We discussed the noisiness of the visualization when there are many documents present, which exposed the need for users to focus on a single document when cleaning the data. While collocations are a common tool for studying text, in this particular task the question of which entities appear in the same sentence is of relatively limited importance. We decided to prioritize edges between documents, mentions, and entities while initially hiding the edges denoting collocations.
The final evaluation presented us with a new list of requirements and feature requests, which are the subject of future work. The need for undo/redo actions and action history was expressed alongside the wish to see specific mentions in the text context. Finally, the need to merge recognized entity nodes arose, due to errors in named entity linking process.

### v2

@@ Abstract
We present an interactive visualization system for exploring named entities and their relationships across document collections. The system is designed around a graph-based representation that integrates three types of nodes: documents, entity mentions, and entities. Connections capture two key relationship types: (i) identical entities across contexts, and (ii) co-locations of mentions within documents. Multiple coordinated views enable users to examine entity occurrences, discover clusters of related mentions, and explore higher-level entity group relationships. To support flexible and iterative exploration, the interface offers fuzzy views with approximate connections, as well as tools for interactively editing the graph by adding or removing links, entities, and mentions, as well as editing entity terms. Additional interaction features include filtering, mini-map navigation, and export options to JSON or image formats for downstream analysis and reporting. This approach contributes to human-centered exploration of entity-rich text data by combining graph visualization, interactive refinement, and adaptable perspectives on relationships. The deployment of the system is available at: https://ul-fri-lgm.github.io/NERVIS/.
@@ 4 System Design
NERVIS is an interactive web-based system for graph visualization and editing, implemented using HTML, CSS, TypeScript, and React framework [^1] https://react.dev. The implementation of the system is available at https://github.com/UL-FRI-LGM/NERVIS. The system is designed around a modular architecture that separates data management, visualization, and user interaction. At its core, NERVIS maintains structured data models that represent nodes, edges, and associated metadata, enabling efficient storage, retrieval, and manipulation of graph information. These models feed into a graph visualization pipeline, which computes layouts, applies styling and interaction rules, and renders the graph for exploration and analysis.
@@ 4.2 Visualization Pipeline
Figure 2: NERVIS visualization pipeline. Orange and gray nodes indicate interactable and uninteractable stages, respectively. Green nodes indicate intermediary data representations.
@@ 5 Expert Evaluation
After initial consultation set the stage and expectations, we performed two rounds of expert system evaluation with synthetic data. After preliminary evaluation, the feature requests were prioritized, and the most important ones were implemented. After the final round, the expert judged the system to be mature enough for use in a real-life setting. They also composed the list of possible future improvements, additions, and extensions.
In the initial consultation, we discussed software goals and previous experience when handling named-entity data with existing software. One of the key things pointed out during this consultation was the burden of human labor when dealing with computer-generated labels. For historians, a recall-focused approach is preferred to an accuracy-focused one. They prefer broad sets of result candidates, which they can browse through and edit manually as opposed to only seeing a single most likely result. This encouraged the design choice to develop a tool where the deletion of nodes and edges is the most common user operation, as opposed to creating new ones. A high recall approach inevitably generates relatively high amounts of noise, which we planned to tackle with a variety of content filters that allow researchers to focus on specific documents, entities, or types when cleaning the data. Since the users of Digital Humanities tools often do not have an engineering background and often struggle with the installation process, we decided on the web platform, which also makes the system broadly available.
After the prototype was tested, we followed up with our second consultation. At this stage, the tool received generally positive feedback, with feature requests focusing on the most common user actions. As already mentioned, with high recall results, most user actions involve deleting nodes and edges, which was made additionally accessible with the implementation of hotkeys for quick node/edge deletion. We discussed the noisiness of the visualization when there are many documents present, which exposed the need for users to focus on a single document when cleaning the data. While collocations are a common tool for studying text, in this particular task, the question of which entities appear in the same sentence is of relatively limited importance. We decided to prioritize edges between documents, mentions, and entities while initially hiding the edges denoting collocations.
The final evaluation presented us with a new list of requirements and feature requests, which are the subject of future work. The need for undo/redo actions and action history was expressed alongside the wish to see specific mentions in the text context. Finally, the need to merge recognized entity nodes arose due to errors in the named entity linking process.